A body is given +1C charge. How many electrons added or removed from the body?
1
Expert's answer
2019-06-25T09:10:09-0400

We can find how many electrons added or removed from the body from the formula:


"Q = Ne,"

here, "Q = +1C" is the charge of the body, "e = 1.6 \cdot 10^{-19} C" is the charge of the electron, "N" is the number of electrons.

Then, we get:


"N = \dfrac{Q}{e} = \dfrac{1 C}{1.6 \cdot 10^{-19} C} = 6.25 \cdot 10^{18} electrons."

Answer:

"N = 6.25 \cdot 10^{18} electrons."